Consider the system of equations shown below.

x - 4y + 5z = 8

-7x + 14y + 4z = -28

3x - 6y + z = 12

If the system is consistent, then write the solution in the form x = xp + xh, where xp is a particular solution of Ax = b and xh is a solution of Ax = 0. (If the system is inconsistent, enter INCONSISTENT in both matrices.)

X = [ ] + t [ ]
Answer #1

Given n-uy + 5z = 8 -72+144+42=-28 30 - 64+ z = 12 Augmented matrix is au 5187 1-7144-28 (3 -6 11 12 R2 R2 +7R R3 R3-3R sau 5an an. take -lung +390 =0 lung= 39.03. = 39 dB and di-uda +5 as=0 diru ( 3 ) ene 1593 to >> 01 - 78 dB + 50370 5 ml- up az o